The FBI wants to determine the effectiveness of their 1010 Most Wanted list. To do so, they need to find out the fraction of people who appear on the list that are actually caught.

Step 1 of 2 :  

Suppose a sample of 791791 suspected criminals is drawn. Of these people, 546546 were not captured. Using the data, estimate the proportion of people who were caught after being on the 1010 Most Wanted list. Enter your answer as a fraction or a decimal number rounded to three decimal places.
Answer #1

Proportion who got caught = (791 - 546)/791 = 35/113 = 0.310